Session: May 2004 (session 2)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (C. Bortolotti at Fri Jun 4 04:41:41 2004) Minor failures : RFI
VC01 and VC02 frequency is out of IF BW, RFI on VC03 and VC04.

Torun (K. Borkowski at Fri Jun 4 05:40:07 2004) Minor failures : RECO
Unstable tape motion due to a problem with internal synchronizaton of the recorder. No other known problems.

Effelsberg (A. Kraus at Fri Jun 4 09:17:06 2004) Success :

Onsala (Michael Lindqvist at Fri Jun 4 09:35:32 2004) Success :
Success OR Failure? During the previous experiment (N04L2) we lost phasecal. However, it could also be a problem with the 5MHz distribution which means that the system became unlocked. After the end of EG030 we turned phasecal on and could see it directly! Thus, we can't say if we had a problem with 5 MHz distribution (which means Failure) OR the phasecal (which means Sucess) during EG030. Since phasecal should be switch of during the experiment we can't say when it came back.

Westerbork (Tony Foley at Fri Jun 4 14:50:54 2004) Success :
Wb used in phased array mode. No known problems

Jodrell_Bank (Alastair Gunn at Sat Jun 5 04:47:01 2004) Success :
MkII was used. No known problems.

Cambridge (Alastair Gunn at Sat Jun 5 04:48:05 2004) Minor failures :
Telescope control fault between 02:04 and 02:08 UT. Possibly 1 or 2 scans lost. Signal only present in BBC01 and BBC02, as requested. No other known problems.

Arecibo (Tapasi Ghosh at Sat Jun 5 14:14:38 2004) Success :
No problem at this end

Noto (P. Cassaro at Tue Jun 8 09:52:39 2004) Success :
No known problems